# Constrained dynamics and confinement in the two-dimensional quantum Ising model

### v2

@@ Abstract
We investigate the dynamics of the quantum Ising model on two-dimensional square lattices up to $16\times 16$ spins. In the ordered phase, the model is predicted to exhibit dynamically constrained dynamics, leading to confinement of elementary excitations and slow thermalization. After demonstrating the signatures of confinement, we probe the dynamics of interfaces in the constrained regime through sudden quenches of product states with domains of opposite magnetization. We find that the nature of excitations can be captured by perturbation theory throughout the confining regime, and identify the crossover to the deconfining regime. We systematically explore the effect of the transverse field on the modes propagating along flat interfaces and investigate the crossover from resonant to diffusive melting of a square of flipped spins embedded in a larger lattice.
@@
One possible reason for slow thermalization is the emergence of dynamical constraints which approximately restrict the time evolution to certain subspaces of the full Hilbert space. A prominent example are systems with confinement [7], where pairs of elementary excitations experience attraction which increases with increasing separation, and thus form long-lived bound states. This is, for example, the reason why quarks at low temperatures do not exist as isolated particles, but are bound in mesons or hadrons.
An analogous mechanism is also found in Ising spin chains [8]. At small transverse fields, the low energy excitations are domain walls, and pairs of them feel a confining force provided by the longitudinal field [9, 10, 11, 12, 13] or interactions beyond nearest neighbour [14]. The direct correspondence between confined domain walls in spin systems and pairs of charges confined by a gauge field in lattice gauge theories makes the Ising chain a useful toy model for studying thermalization of bound states via meson production [15, 16, 17, 18]. Domains of spins anti-aligned with the longitudinal field are bubbles of false vacuum, and probing their dynamics is crucial for understanding false vacuum decay [19, 20].
In two dimensions (2D), the confining potential comes from the lattice itself. An excitation has to flip neighbouring spins to propagate, thus adding domain walls. At small transverse fields such processes are very energetically costly, and thus strongly suppressed. Confinement has initially been studied in ladders [21, 22], by orthogonally coupling many chains described by the Ising field theory in the scaling limit [23, 7] and recently in a 2D lattice with infinite tensor network methods [24, 25]. These approaches describe the system in the thermodynamic limit and hence cannot probe dynamics of inhomogeneous initial states and domain walls.
In this paper, we use tree tensor networks (TTNs) to explore the constrained dynamics of the quantum Ising model on 2D square lattices of up to $16\times 16$ spins, through quenches of product states to finite transverse fields. Starting with the completely polarized state, we demonstrate signatures of confinement and investigate the nature of low-energy excitations that drive the dynamics. Then, we turn our attention to two representative types of domain walls, a flat interface between two spin species and a square domain embedded in a larger lattice, to explore the melting of interfaces in the confining regime. We characterize the surface modes that are known to dominate the dynamics in the limit of small transverse field [26, 27], and investigate the crossover towards diffusive melting. We find that the crossover corresponds to the dynamical phase transition from the ordered to disordered phase and we characterize the nature of domain melting in both regimes. Finally, we discuss the feasibility of implementing our numerical calculations on quantum simulators and suggest an outlook of possible extensions of the presented work.
Model and methods - The Hamiltonian of the two-dimensional quantum Ising model is
@@
with $x$- and $z$-Pauli matrices $\sigma^{x}$ and $\sigma^{z}$, $J$ the interaction, and $g$ the transverse field strength. Sum indices run over sites of a $N\times N$ square lattice, with $\langle i,j\rangle$ indicating a sum over nearest neighbours. We choose to work with periodic boundary conditions throughout, but the physics observed here does not depend on this choice.
Throughout the paper, we investigate instantaneous quenches of product states with spins aligned in the $z$-direction (eigenstates at $g=0$) to finite values of $g$. We encode them into tree tensor networks [28, 29, 30] and propagated in real time using the time-dependent variational principle (TDVP) [31, 32], with a timestep of $0.005J$. Our results are generally sufficiently converged for the maximal bond dimension $\chi=100$. See Supplemental Material (SM) for technical details and convergence analysis.
Broadly speaking, the accuracy of tensor network methods is bound by entanglement. After instantaneous quenches of product states it typically grows, at worst linearly with time [33]. This means that the resources required for accurate tensor network calculations grow exponentially. However, as shown in this work and elsewhere [9, 7, 24], the physical regime investigated here is noteworthy precisely because entanglement grows slowly in time. This makes TTNs a particularly apt tool to attack the present problem.
@@
Figure 1: Confinement in the 2D quantum Ising model. (a) Short time dependence of magnetization $\left\langle\sigma^{z}(t)\right\rangle$ for a set of $g/J$. (b) $\left\langle\sigma^{z}(t)\right\rangle$ for longer times and small $g/J$. (c) Time dependence of entanglement entropy $S$ for a bipartition of two neighbouring spins and the rest of the system. $S_{\mathrm{max}}$ is the maximal entanglement entropy of the bipartition (in this case $S_{\mathrm{max}}=2\ln 2$). (d) Time dependence of a horizontal cut of the connected correlation function $C_{ij}$ where $i=4$ for a set of $g/J$.
@@
The time evolution of the entanglement entropy $S$ is shown in Fig. 1(c). We choose a bipartition between two neighbouring spins and the rest of the system, but all other bipartitions produce similar results. At small $g/J$, we observe an oscillatory pattern with the system periodically returning very close to a product state with a frequency of approximately $\frac{\pi}{2}J$. Interestingly, the frequency is half the frequency of the oscillations of $\left\langle\sigma_{z}\right\rangle$. The oscillations are washed out as $g$ is increased and the density of excitations becomes large enough that they interact, leading to thermalization [35].
The spin-spin correlations also exhibit signatures of confinement, see Fig. 1(d), where we plot a vertical cut of $C_{ij}$ with spin $i$ chosen in the fourth row. The spread of correlations is strongly suppressed for small $g/J$. At $g/J=0.5$, we find non-monotonous behaviour with the same frequency as the oscillations of entanglement. The light-cone component becomes more apparent as $g$ is increased, and at $g/J=1.5$ it is clearly the dominant contribution. It should be mentioned that the light-cone never vanishes completely, but is only suppressed. Its apparent disappearance is an artefact of the choice of a plotting scale [9, 14].
@@
Nature of excitations - The oscillating behaviour of local observables implies that the dynamics is generated by repeated creation and annihilation of local excitations. Their nature can be inferred from the spectral density of local observables [9, 35, 14, 24]. In Fig. 2 we show the results of the Fourier transform of the magnetization from Fig. 1(b) between $tJ=10$ and $tJ=80$, denoted by $\mathcal{F}(\sigma^{z})$. We multiply the timeseries by the Hamming window to exclude possible transient events at the edges of the interval [36, 37]. An example of the spectral density for two values of $g/J$ is shown in Fig. 2(a), while a heatmap for a broader range is in Fig. 2(b).
The three peaks match the excitation energies of the dominant dynamical processes. The most prominent peak, starting at $\omega=8J$ for $g=0$, corresponds to the transition between the initial state and a state with one flipped spin, $\psi_{1}$. The peak at $12J$ corresponds to a transition between $\psi_{0}$ and $\psi_{2}$, a state where two neighbouring spins are flipped, with a total domain wall length of $6$. Finally, the peak at $4J$ is produced by the transition between $\psi_{1}$ and $\psi_{2}$.
The $g$-dependence of the excitation energies can be captured by perturbative corrections to the product states. The expressions up to the second order in $g$ are:
@@
Note that $N$ is the linear size of our system, which consists of $N^{2}$ sites in total. For complete calculations see the SM.
@@
The black dashed lines in Fig. 2(b) indicate the transition energies obtained as differences of these energies, denoted $\Delta_{ij}=E_{j}-E_{i}$. They match the positions of the spectral peaks to surprisingly large values of transverse field, up to $g\sim J$. This is somewhat surprising for expressions which are technically only accurate up to the second order in $g$, and indicates that the contribution of the higher excited modes is negligible throughout the confining regime.
Note that aside from the extensive correction common to all states, we do not find a second-order contribution to $E_{2}$. An intuitive interpretation is that in second order the corrections that come from coupling $\psi_{2}$ to $\psi_{1}$, and the ones coming from $\psi_{3}$ (superposition of connected three-spin clusters, total domain length of 8), originate from the same physical process, ie. adding or removing a spin from an existing cluster. Both add two domain walls, and being symmetric in this sense, they cancel out. The processes that couple $\psi_{2}$ to $\psi_{0}$ only appear at fourth order, and we expect that accounting for those would reproduce the curvature of the right and left peaks in Fig. 2(b).
The spectrum of excitations hints at the underlying reasons for quick thermalization beyond $g/J\sim 2$. The $\Delta_{12}$ peak indicates the cost of flipping a spin next to an already existing spin flip – growing a spin cluster by one. Quenching to the point where this process is energetically as expensive as creating a single spin flip populates not only the first few but also higher excited states, resulting in the formation of larger clusters and thus thermalization on a timescale of $1/J$.
Because the dynamics of confined systems is dominated by low energy excitations, it can be accurately described by a simple effective model [16, 14, 24]. This is typically done by projecting the Hamiltonian to a subspace of low-energy excitations (equivalents of our $\psi_{0}$, $\psi_{1}$ and $\psi_{2}$). However, this approach leads to certain inconsistencies. Namely, the matrix element between the ground and first excited state is extensive (indeed, in our case we find $\left\langle\psi_{0}\right|H\left|\psi_{1}\right\rangle=gN$), while others are independent of $N$. Therefore, an effective model including $\psi_{0}$ predicts a spectral gap (equivalent of $\Delta_{01})$ that diverges with system size. To accurately reproduce the excitation energies, the initial vacuum state (equivalent of $\psi_{0}$) should be excluded from the low energy model, and its mass set exactly to zero.
Perturbative calculations imply that the origin of the problem is in the arbitrary truncation of the subspace of the effective model (see SM for details). We find that the extensive corrections come from a state coupling to an extensive number of excitations with one more and one fewer spin flip – it is always possible to flip a decoupled spin in approximately $N^{2}$ places for an energy penalty of $8J$. The contributions coming from the state with one more and one less spin flip cancel out in the perturbative calculations, and we only find an overall $g^{2}N^{2}/8J$ correction to all states, so that energy differences remain independent of system size. However, if the Hilbert space is first truncated into an effective model, the coupling between the last retained and the first truncated state is not taken into account, the extensive contributions do not cancel out, and the gaps grow with system size. While the two approaches produce a similar energy diagram, an advantage of the perturbative approach is that it can also provide corrections to the eigenstates.
@@
Interfaces - Next, we probe the dynamics of interfaces by investigating quenches of inhomogeneous initial states. We consider two examples: the stripe, an initial state where two domains of opposite polarized spins are separated by a flat interface, and the square, where the flipped spins are arranged into a square domain.
@@
Confinement importantly affects interface dynamics. Whereas interfaces are typically expected to diffusively melt, in the confining phase they tend to remain stable up to very long times. This is because classically resonant processes dominate the dynamics for small $g/J$, ie. only spins with two neighbours of each spin species can flip, as the initial and final state have the same energy. These processes only locally reshape the domain wall, and thus its total length emerges as a dynamic constraint. The interface dynamics is well understood in the limit of infinite $J$ where non-resonant processes are explicitly forbidden [26, 27]. Slow thermalization comes as a consequence of the subsequent fragmentation of Hilbert space into sectors labeled by the domain wall length, and further into so-called dynamically decoupled Krylov subspaces; two domains cannot merge through local reshaping of the interface if they are separated by multiple lattice sites [38, 39].
@@
Figure 3: Spread of correlations near an interface. (a) Sketch of the initial state, with blue representing $\uparrow$ and red $\downarrow$ spins. Black lines indicate the cuts shown in (d). (b) Sketch of the resonant process along the interface. (c) Heatmap of the spectral density of $\left\langle\sigma^{z}\right\rangle(t)$ for a spin at the interface for a range of $g/J$. White dashed lines correspond to the transition energies of a freely propagating edge mode. (d) Horizontal ($x$) and vertical ($y$) cuts of the connected correlations function $C_{ij}$ with respect to the spin at $(4,4)$ (white square in (a)). Blue dashed lines are ±$\pm 4gt$, showing that the interface mode carries the correlations.
@@
In this case, the lowest excitation above the initial state is a spin flip at the interface. It creates two additional domain walls and costs $2\times 2J$. However, once this excitation is present, flipping the neighbouring interface spin does not come with any additional energy penalty, see Fig. 3(b). Such excitations are thus allowed to freely propagate along the boundary, with the expected dispersion $E(k)=4J\left(1+g\cos(k)\right)$ [40].
@@
The spectral density of $\left\langle\sigma^{z}\right\rangle$ for a spin at the interface is shown in Fig. 3(b). The most prominent feature is the excitation which starts at $\omega=4J$ for $g=0$ and fans out into a continuum with increasing $g$, as expected of a freely propagating mode. White dashed lines correspond to the expected envelopes of the transition energies $\omega=4J\pm 4g$. The continuum is split into discrete peaks as a consequence of the finite size of the lattice.
@@
Vertical and horizontal cuts of the connected correlation functions are shown in Fig. 3(c). Because of the interface, the spread of correlations is not the same in the horizontal and vertical directions. In the horizontal direction (orthogonal to the interface) correlations show signatures of confinement equivalent to results in Fig. 1. Interestingly, the correlation spread is symmetric, meaning that the confined excitations spread across the interface into the domain with opposite polarisation in the same way as within the domain. On the other hand, the vertical (y) cuts show very different behaviour. The freely propagating interface mode spreads the correlations along the interface with a velocity proportional to $g$ (blue dashed lines in Fig. 3(c)), even for $g\ll J$. As $g$ is further increased the dynamics lose their $x$-$y$ asymmetric nature, and at large $g$ correlations spread equally in either direction.
@@
Finally, we consider the melting of an $8\times 8$ square of $\uparrow$ spins embedded in a $16\times 16$ lattice and surrounded by $\downarrow$ spins. Fig. 4(a) shows snapshots of magnetization for small and intermediate $g/J$, while Fig. 4(b) shows the time dependence of magnetization for a set of spins close to a corner of the square. For extended plots of magnetization as well as correlation spread, see the SM.
@@
As shown in Refs. [26, 27], in the $J\rightarrow\infty$ limit and for an infinitely large corner the melting process can be understood as a series of resonant spin-flip processes. Because the corner spin has two neighbours of each spin species it flips on a timescale set by $1/g$. After the corner spin is flipped, its two neighbours can resonantly flip as well. This triggers a propagating edge mode in the same way as in Fig. 3. The process is illustrated in the magnetization snapshots for $g/J=0.1$ in Fig. 4(a), and results in the square melting into a cross-like shape.
@@
A more quantitative view is presented in Figs. 4(b,c). At small $g/J$ the edge mode dominates. For spins along the interface shown in panel (c), the delay of melting is thus linear in the distance from the corner. Probing deeper into the square along the diagonal (Fig. 4(b)), we find that spins near the center stay inert up to relatively long times. Note that here we rescale time with $g$, and find that the small $g/J$ curves in Fig. 4(b,c) follow the same trend. This indicates that the smaller energy scale $g$ dominantly drives the dynamics.
@@
Increasing $g$ causes deviations from this simple picture and a crossover towards a regime where all spins at the interface melt simultaneously. The amplitude of magnetization oscillations is gradually reduced, and at $g/J=0.9$ we find a monotonous decrease of $\left\langle\sigma^{z}\right\rangle$ for all spins of the square simultaneously. The melting produces waves of magnetization radiating from sides of the square into the environment (see bottom panel of Fig. 4(c), and SM extended plots). This is a sign that the dynamics is no longer confined, and that the length of the domain wall is not conserved.
@@
Realization on a quantum simulator - The Ising model is naturally realized in various quantum simulation platforms. However, being two-dimensional, the presented setup is most directly applicable to neutral atom arrays and within reach of existing experiments. More concretely, Ref. [41] demonstrates control over 2D arrays of hundreds of Rydberg atoms realizing a nearest-neighbour antiferromagnetic Ising model with the interaction on the order of 222\,2MHz and coherence time of $20\mu$s. When driving the transition between the two spin states at around 111\,1Mhz, the timescales shown in Fig. 4 correspond to experimental times of a few microseconds. If one chose to relax the nearest neighbour constraint the atoms could be moved much closer together, leading to stronger interactions and thus much faster dynamics. See for example Refs. [42, 43] for experimental and Ref. [44] for a similar theoretical proposal.
@@
Outlook - We used tree tensor networks to study constrained dynamics in the ordered phase of the two-dimensional quantum Ising model. We showed the signatures of confinement of quasiparticles and extracted the spectrum of excitations from oscillations of the magnetization. The spectrum can be adequately reproduced by perturbation theory up to $g\sim J$, while the failure at larger $g$ is related to thermalization. Then, we investigated the melting of interfaces of domains with opposite spins, how it depends on confinement and how it changes with increasing transverse field. We find a crossover between the constrained dynamics of classically resonant processes in the limit of small $g$ and diffusive melting at large $g$.
@@
We see various other open questions that could be addressed either on a quantum simulator, or with minimal alterations of our setup. One interesting avenue is the exploration of Floquet dynamics through quenches with a time-dependent transverse field. Because relatively long timescales are reachable, it is possible to investigate the regime of small and intermediate frequencies, and thus explore the regime beyond the $1/\omega$ corrections of the Magnus expansion. Floquet dynamics have recently been studied in a related system, where a discrete time crystal was identified in the prethermal state [46].
Another option is to investigate false vacuum decay. In two dimensions, dynamical confinement and the presence of false vacuum are decoupled – confinement is provided by the lattice, while the false vacuum is invoked by a longitudinal field. A starting point would be to investigate how the melting in Fig. 4 depends on longitudinal field, which makes the square either a false or a true vacuum domain.
Expanding on this idea, it would be interesting to probe the scattering of false vacuum bubbles. Such calculations have been performed in one dimension [47], but many more scattering channels are expected to arise in two dimensions. However, to clearly discern scattering from lattice effects one would have to significantly increase the system size, possibly beyond the reach of classical simulations.

### v3

@@ Abstract
We investigate the dynamics of the quantum Ising model on two-dimensional square lattices up to 16 × 16 spins. In the ordered phase, the model is exhibits dynamically constrained dynamics, leading to confinement of elementary excitations and slow thermalization. After demonstrating the signatures of confinement, we probe the dynamics of interfaces in the constrained regime through sudden quenches of product states with domains of opposite magnetization. We find that the nature of excitations can be captured by perturbation theory throughout the confining regime, and identify the crossover to the deconfining regime. We systematically explore the effect of the transverse field on the modes propagating along flat interfaces and investigate the crossover from resonant to diffusive melting of a square of flipped spins embedded in a larger lattice.
@@
One possible reason for slow thermalization is the emergence of dynamical constraints which approximately restrict the time evolution to certain subspaces of the Hilbert space. A prominent example are systems with confinement [7], where pairs of elementary excitations experience attraction which increases with increasing separation, and thus form long-lived bound states. The resulting slow thermalization has been extensively studied in Ising spin chains. At small transverse fields, the low energy excitations are domain walls, and pairs of them feel a confining force provided by the longitudinal field [8, 9, 10, 11, 12, 13] or long-range interactions [14].
In the two-dimensional (2D) quantum Ising model, the dynamical constraint comes from the lattice itself. An excitation has to flip neighbouring spins to propagate, thus creating domain walls. At small transverse fields such processes are energetically costly, and thus strongly suppressed. Confinement has initially been studied in ladders [21, 22], by orthogonally coupling chains described by the Ising field theory in the scaling limit [23, 7] and with infinite tensor network methods [24, 25].
In this letter, we use tree tensor networks (TTNs) to explore the constrained dynamics of the quantum Ising model on 2D square lattices of up to $16\times 16$ spins, through quenches of product states to finite transverse fields. We demonstrate signatures of confinement for a completely polarized state and investigate the nature of low-energy excitations that drive the dynamics. Then we turn our attention to inhomogeneous states with domain walls. We consider two representative types: a flat interface, and a square domain embedded in a larger lattice.
In the limit of infinite interaction strength, the dynamics of such states is dominated by surface modes which are confined to and propagate along domain walls [26, 27]. We investigate the transition towards diffusive melting and coarsening of interfaces at finite transverse fields, and characterize the nature of domain melting in intermediate regimes. We conclude by discussing the feasibility of implementing our calculations on quantum simulators and suggest a number of possible extensions of the presented work.
Model and methods — The Hamiltonian of the two-dimensional quantum Ising model is
@@
with $x$- and $z$-Pauli matrices $\sigma^{x}$ and $\sigma^{z}$, $J$ the interaction, and $g$ the transverse field strength. Sum indices run over sites of a $N\times N$ square lattice, with $\langle i,j\rangle$ indicating a sum over nearest neighbours. We work with periodic boundary conditions throughout, but the observed physics does not depend on this choice.
Throughout, we investigate instantaneous quenches of product states with spins aligned in the $z$-direction (eigenstates at $g=0$) to finite values of $g$. We encode them into tree tensor networks [28, 29, 30] and propagate in real time using the time-dependent variational principle (TDVP) [31, 32], with a timestep of 0.005 $0.005\ 1/J$. Our results are generally sufficiently converged for the maximal bond dimension $\chi=100$. See Supplemental Material (SM) for technical details and convergence analysis.
@@
Figure 1: Confinement in the 2D quantum Ising model. (a) Short time dependence of magnetization $\left\langle\sigma^{z}(t)\right\rangle$ for a set of $g/J$. (b) $\left\langle\sigma^{z}(t)\right\rangle$ for longer times and small $g/J$. (c) Entanglement entropy $S$ for a bipartition of two neighbouring spins and the rest of the system. $S_{\mathrm{max}}$ is the maximal entanglement entropy of the bipartition (in this case $S_{\mathrm{max}}=2\ln 2$). (d) Horizontal cuts of $C_{ij}$ through the middle of the system.
@@
The time evolution of the entanglement entropy $S$ is shown in Fig. 1(c). We choose a bipartition between two neighbouring spins and the rest of the system, but all other bipartitions produce similar results. At small $g/J$, we observe the system periodically returns close to a product state with a frequency of approximately $\frac{\pi}{2}J$, half of the frequency of $\left\langle\sigma_{z}\right\rangle$. The oscillations are washed out as $g$ increases and the density of excitations becomes large enough to interact, leading to thermalization [35].
Fig. 1(d) shows vertical cuts of $C_{ij}$ with spin $i$ chosen in the fourth row. The spread of correlations is strongly suppressed for small $g/J$. At $g/J=0.5$, we find non-monotonous behaviour with the same frequency as the oscillations of entanglement. The light-cone component becomes more apparent as $g$ is increased, and at $g/J=1.5$ it is clearly the dominant contribution. It should be mentioned that the light-cone never vanishes completely, but is only suppressed. Its apparent disappearance is an artefact of the choice of a plotting scale [9, 14].
@@
Nature of excitations — The oscillating behaviour of local observables implies that the dynamics is generated by creation and annihilation of local excitations. Their nature can be inferred from the spectral density of local observables [9, 35, 14, 24]. In Fig. 2 we show the results of the Fourier transform of the magnetization from Fig. 1(b) between $tJ=10$ and $tJ=60$. Altering the upper bound of the interval between $tJ=30$ and $tJ=80$ does not change the spectra considerably. We apply the Hamming window to the data to exclude possible transient events at the edges of the interval [36, 37]. An example of the obtained spectral density is shown in Fig. 2(a). A heatmap for a broader range is in Fig. 2(b).
The three peaks correspond to excitation energies of the dominant processes. The most prominent one, at $\omega\approx 8J$, corresponds to the transition between the initial state and a state with one flipped spin, $\psi_{1}$. The peak at $12J$ corresponds to a transition between $\psi_{0}$ and $\psi_{2}$, a state where two neighbouring spins are flipped, total domain wall length of $6$. Finally, the peak at $4J$ corresponds to the transition between $\psi_{1}$ and $\psi_{2}$.
The transition energies are captured by perturbative corrections to the product states. Up to the second order in $g$ we find:
@@
Note that $N$ is the linear size of our system, consisting of $N^{2}$ sites in total.
@@
The black dashed lines in Fig. 2(b) indicate the transition energies, $\Delta_{ij}=E_{j}-E_{i}$. They match the positions of the spectral peaks to surprisingly large values of transverse field, up to $g\sim J$. This agreement is surprisingly good for expressions of second order in $g$, and indicates that the contribution of the higher excited modes is negligible throughout the confining regime.
See the End Matter section for a deeper interpretation of the obtained expressions, and the SM for complete calculations.
@@
Interfaces — Next, we probe the dynamics of interfaces by investigating quenches of inhomogeneous initial states. We consider two initial states: the stripe, where two domains are separated by a flat interface, and the square, where the flipped spins are arranged into a square domain.
@@
The presence of dynamical constraints importantly affects interface dynamics. Whereas interfaces are typically expected to diffusively melt, in the confining phase they tend to remain stable up to very long times. This is because classically resonant processes dominate the dynamics for small $g/J$, ie. only spins with two neighbours of each spin species can flip, as the initial and final state have the same energy. These processes only reshape the domain wall, and thus its total length emerges as a dynamical constraint. The interface dynamics is well understood in the limit of infinite $J$, where non-resonant processes are explicitly forbidden [26, 27]. Slow thermalization comes as a consequence of the subsequent fragmentation of Hilbert space into sectors labeled by the domain wall length, and further into so-called dynamically decoupled Krylov subspaces; two domains cannot merge through local interface reshaping if they are spatially separated [38, 39].
@@
Figure 3: Spread of correlations near an interface. (a) Sketch of the initial state, with blue representing $\uparrow$ and red $\downarrow$ spins. Black lines indicate the cuts shown in (d). (b) Sketch of the resonant process along the interface. (c) Heatmap of the spectral density of $\left\langle\sigma^{z}\right\rangle(t)$ for a spin at the interface for a range of $g/J$. White dashed lines correspond to the transition energies of a freely propagating edge mode. (d) Horizontal ($x$) and vertical ($y$) cuts of the connected correlations function $C_{ij}$ with respect to the spin at $(4,4)$ (white square in (a)). Blue dashed lines are ±$\pm 4gt$, showing that the interface mode carries the correlations. (e) Cuts of magnetization along a row in a $16\times 8$ system, with periodic boundary conditions along the shorter, and open along the longer dimension. The system is initialized with two polarized domains and an interface between spins $7$ and $8$ (black dashed line). The cuts are taken at evenly spaced times between $tJ=0$ (darkest brown) and $tJ=5$ (lightest yellow).
@@
The lowest excitation above the stripe is a spin flip at the interface. It creates two additional domain walls and costs $2\times 2J$. Once this excitation is present, flipping the neighbouring interface spin does not come with any additional energy penalty, see Fig. 3(b). Such excitations are thus allowed to freely propagate along the boundary, with the expected dispersion $E(k)=4J\left(1+g\cos(k)\right)$ [40].
@@
The spectral density of $\left\langle\sigma^{z}\right\rangle$ for a spin at the interface is shown in Fig. 3(c). The most prominent feature is at $\omega=4J$ for $g=0$ and fans out into a continuum with increasing $g$, as expected of a freely propagating mode. White dashed lines correspond to the expected envelopes of the transition energies $\omega=4J\pm 4g$. The continuum is split into discrete peaks as a consequence of the finite size of the lattice.
@@
Vertical and horizontal cuts of the connected correlation functions are shown in Fig. 3(d). Because of the interface, the spread of correlations is not the same in the horizontal (x) and vertical (y) directions. In the $x$ direction (orthogonal to the interface) correlations show signatures of confinement equivalent to Fig. 1. Interestingly, the correlation spread is symmetric, meaning that the confined excitations spread across the interface into the domain with opposite polarisation in the same way as within the domain. On the other hand, the $y$ cuts show very different behaviour. The freely propagating interface mode spreads the correlations along the interface with a velocity proportional to $g$ (blue dashed lines in Fig. 3(d)), even for $g\ll J$. As $g$ is further increased, the correlations gradually lose their $x$-$y$ asymmetric nature, and at large $g$ spread equally in either direction.
Due to the scale separation between spin-flips at the interface and in the bulk the crossover from constrained to thermalizing dynamics exhibits an intermediate regime, where the interface broadens while the bulk remains inert. While it only faintly appears in the correlations, the regime is clearly visible in the magnetization cuts, Fig. 3(e). Here we use a $16\times 8$ system on a cylinder (periodic boundary conditions along the shorter and open along the longer dimension).
The three regimes are discerned by the width of the active area. Deep in the confining regime (see $g/J=0.3$) only the interface spins are active through the edge mode mechanism. At large $g$ ($g/J=2.5$), the interface does not present an energetic barrier, and the thermalization occurs indiscriminately at the interface and in the bulk. In the intermediate regime ($g/J\sim 1-2$), we find that the interface melts and broadens, while the bulk remains inert. This results in a gradient of magnetization between the two domains at late times. A larger set of cuts in the intermediate regime is provided in the SM.
The presence of multiple regimes affects the course of thermalization and the final thermal state. Each process has its associated energy scale, and thus a corresponding thermalizing time scale. In such cases, the thermalization is expected to occur in stages, as the system relaxes to the Gibbs ensemble corresponding to the subspace of the Hilbert space accessible at a given timescale [Birnkammer2022]. This can affect the nature of the final thermal state [Zadnik2023]. The influence of the domain wall shape on the thermalization remains an interesting open question.
@@
Finally, we consider the melting of an $8\times 8$ square domain embedded in a $16\times 16$ lattice. Fig. 4(a) shows the evolution of the magnetization for small and intermediate $g/J$, while Fig. 4(b) shows the time dependence of the magnetization for a set of spins close to a corner of the square. For extended plots of the magnetization as well as the correlation spread, see the SM.
@@
In the $J\rightarrow\infty$ limit and for an infinitely large corner, the melting process can be understood as a series of resonant spin-flip processes [26, 27]. The corner spin has two neighbours of each spin species, so it flips on a timescale set by $1/g$. After the it is flipped, its two neighbours resonantly flip as well. This triggers a propagating edge mode in the same way as in Fig. 3. The process results in the square melting into a cross-like shape, as shown in Fig. 4(a) for $g/J=0.1$.
@@
A more quantitative view is presented in Figs. 4(b,c). Due to the dominating edge mode at small, $g/J$ the spins along the interface (Fig. 4 (c)) melt with a delay which is linear in their distance from the corner. Probing deeper into the square along the diagonal (Fig. 4(b)), we find that spins near the center stay inert up to relatively long times.
@@
Increasing $g$ causes deviations from this simple picture and a crossover towards a regime where all spins at the interface melt simultaneously. The amplitude of magnetization oscillations is gradually reduced, and at $g/J=0.9$ we find a monotonous decrease of $\left\langle\sigma^{z}\right\rangle$ for all spins of the square simultaneously. The melting produces waves of magnetization radiating from sides of the square into the environment (see bottom panel of Fig. 4(c)). This is a sign that the dynamics is no longer confined, and the length of the domain wall not conserved.
Realization on a quantum simulator — The Ising model is naturally realized on various experimental platforms. However, being two-dimensional, the presented setup is most directly applicable to neutral atom arrays and within reach of existing experiments.
@@
More concretely, Ref. [41] demonstrates control over 2D arrays of hundreds of Rydberg atoms realizing a nearest-neighbour antiferromagnetic Ising model with the interaction on the order of 222\,2MHz and coherence time of $20\mu$s. When driving the transition between the two spin states at 111\,1Mhz, the timescales shown in Fig. 4 correspond to experimental times of a few microseconds. If one chose to relax the nearest neighbour constraint the atoms could be moved much closer together, leading to stronger interactions and thus much faster dynamics. See for example Refs. [42, 43] for experimental and Ref. [44] for a similar theoretical proposal.
@@
Outlook — We used tree tensor networks to study constrained dynamics in the ordered phase of the 2D quantum Ising model. We showed the signatures of confinement of quasiparticles and extracted the spectrum of excitations from oscillations of the magnetization. The spectrum can be adequately reproduced by perturbation theory up to $g\sim J$. We investigated the melting of interfaces between domains, how it depends on confinement and how it changes with increasing transverse field. We find a crossover between the constrained dynamics of classically resonant processes in the limit of small $g/J$ and diffusive melting at large $g/J$.
@@
Given the demonstrated success of TTNs to accurately simulate the time evolution of dynamically constrained two dimensional systems, we see the potential to tackle various open questions and new avenues of research.
Being very similar in nature to spin systems, the 2D Rydberg arrays in the checkerboard phase also exhibit constrained dynamics [43]. However, the nature of the dynamical constraints, and the properties of basic excitations in such systems are not well understood. Investigation of these properties, as well as their interplay with coarsening of domains are well in reach of the presented method, and subject of future work.
Another promising avenue is the investigation of false vacuum decay in 2D. In 1D, the proposed mechanism for the decay is the nucleation of true vacuum bubbles [19, 20, Vodeb_2024]. Nucleation is driven by the competition between the bulk and interface contributions to the bubble energy. In 1D, the area of the bubble remains constant as it grows, while in higher dimensions it increases with the power of $D-1$. Exploring nucleation in 2 dimensions should thus provide a qualitatively new, and more universally accurate understanding of the dynamics of the false vacuum. A relatively simple starting point would be to investigate how the melting in Fig. 4 depends on longitudinal field, which makes the square either a false or a true vacuum domain.
Expanding on this idea, it would be interesting to probe the scattering of false vacuum bubbles. Such calculations have been performed in one dimension [47], but multiple scattering channels are expected to arise in two dimensions. However, clearly discerning scattering from lattice effects requires a significant increase of the system size, possibly beyond the reach of classical simulations.